OpenStudy (anonymous):

Which function has x-intercepts of 3 and 8 and contains the point (-1, -2)?

OpenStudy (shamil98):

You imply there are options to this problem, can you post them?

OpenStudy (anonymous):

yea one sec

OpenStudy (anonymous):

y= -1/36(x-3)(x-8)

OpenStudy (anonymous):

y= -1/14 (x-3)(x-8)

OpenStudy (anonymous):

y=-1/18(x-3)(x-8)

OpenStudy (anonymous):

y=-1/7(x-3)(x-8)

OpenStudy (anonymous):

thats it

OpenStudy (jdoe0001):

\(\bf x=3\implies x-3=0\implies (x-3)=0\\ \quad \\ x=8\implies x-8=0\implies (x-8)=0\\ \quad \\ \quad \\ y=a(x-3)(x-8)\qquad \textit{if we know that }(-1,-2)\quad x=-1, y=-2\\ \quad \\ \textit{plug those values in, and solve for "a"}\)